OpenStudy (anonymous):

I will fan and medal! If θ is in Quadrant IV and sinθ = -1/2, then cosθ=? the choices are: -(2sqrt(3)/3) -(sqrt(3)/2) sqrt(3)/2 (2sqrt(3)/3)

OpenStudy (crissy15):

@Jacob902

Nnesha (nnesha):

draw a right triangle in quadrant 4  definition of sin \[\large\rm sin = \frac{ opposite }{ hypothounuse }\] cos = adjacent over hypotenuse apply Pythagorean theorem find the 3rd side of the right triangle \[\huge\rm a^2 +b^2 = c^2\]